OpenStudy (anonymous):

Solve the following system of equations: a.) x –2y+z=6 B.) 2x+y –3z= –3 c.) x –3y+3z=10

OpenStudy (anonymous):

solve them for what? you mean put them into a different form such as a.) x=6-2y+z? or a form such as, y= -x - 3 + z??